OpenStudy (anonymous):

What is the value of the expression below after it has been completely simplified? -2/0.25-0.5/+2

OpenStudy (ahsome):

Equation is\[\frac{-2}{0.25}+\frac{-0.5}{2}\] Change so they have the same denominator, here 2 \[\frac{-2*8}{0.25*8}+\frac{-0.5}{2}\] \[\frac{-16}{2}+\frac{-0.5}{2}\] Add common terms \[\frac{-16-0.5}{2}\] \[\frac{-16.5}{2}\] \[-8.25\] Hope that helped, @esther12345678

OpenStudy (ahsome):

Am I right, @iGreen?

OpenStudy (ahsome):

(New to this)